OpenStudy (anonymous):

What is the solution to the equation below? Round your answer to two decimal places. 10 log(4x) = 25 A. x = 316.23 B. x = 93.78 C. x = 79.06 D. x = 1264.91

OpenStudy (wmj259):

1.) Isolate log(4x) by dividing both sides by 10.

OpenStudy (anonymous):

Divide both sides by 10 log 4x = 25/10 = 2.5 Raise 10 to the 2.5 power 4x = 10 ^ (2.5) and then divide the answer by 4 x = 10 ^(2.5) / 4 = 79.056

OpenStudy (anonymous):

log(4x)=5/2

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

@wmj259 I didn't see you there :/ sorry.. but @Jooooj no problem

OpenStudy (wmj259):

@Educated no problem.